i am having trouble with solving this equation: -8/13x=24
Answer by jim_thompson5910(35256)

-8/13x=24 ... Start with the given equation

-8x=24*13 ... Multiply both sides by 13

-8x=312 ... Multiply

x=312/(-8) ... Divide both sides by -8 to isolate x

x = -39 ... Reduce


So the solution is x = -39